• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 16:01:37 on June26,2002

    Entry number 86827

    keyword=Shift summary


    New crew on shift: J.-P. Chen, Kent Paschke, Eric Voutier, Christophe Furget
    ***************

    08:00 Continue data taking on kinematics Q2_k40
    08:51 Target IOC reboot
    08:58 Back in business
    09:33 ROC14 crashed about 20 mn ago without giving us any alarm, care should be taken in the analysis.
    09:36 Back in business
    10:07 DAQ problems, ROC14 does not want to come back...
    10:18 Back in business
    11:12 ROC14 crashes
    11:15 Back in business
    14:37 Beam is a little off is x (-0.083); ask MCC to tune it... does not succeed... then
    forget it!
    14:40 MCC takes beam away to adjust the path length
    15:21 Beam back in the hall

    Run History
    --------

    >- Q2_k40 -<

    1819 .......... Good run / stop to reboot target IOC
    1820 .......... ROC14 crashes
    1821 .......... ROC14 crashes
    1822 .......... junk
    1823 .......... ROC14 crashes
    1824 .......... Good run
    1825 .......... Good run
    1826 .......... Good run
    1827 .......... Good run
    1828 .......... goes to next shift